# Predicting and Understanding Turn-Taking Behavior in Open-Ended Group Activities in Virtual Reality

### v1

@@ 3.3.3. Egocentric Motion Features
Egocentric motion features capture the pose and motion of individual users. In our setup, we extracted egocentric motion features from both the main user and the reference user. For each user, we extracted the position (i.e., x, y, z) and orientation (i.e., pitch, yaw, roll) of a user’s headset and two controllers by both their raw values and velocities. From the raw values and velocities across the 1-second window, we summarized each of the 6 degrees of freedom by their average, minimum, and maximum values. We transformed the features corresponding to the raw values into the body-space coordinate system proposed in Miller et al. (Miller et al., 2023b) and centered the horizontal coordinates (i.e., x, z) such that the individual’s head position is at the origin of the horizontal plane, as opposed to their global coordinates in the virtual space. The rationale behind applying this transformation is so that the raw features are encoding the egocentric poses by mapping the horizontal coordinates and orientations to a coordinate system based on the individual’s forward head direction. This allows the raw features to intuitively encode poses even when individuals move around their physical or virtual environments.
When calculating the velocity of the position and orientations, we derived the velocities from the raw values after transforming the original values into the body-space coordinates without the centering operation. Choosing to not center the coordinates based on the head position allows the features to retain information of the user’s head position velocities in the horizontal plane. When encoding user head yaw (i.e., rotation about the vertical axis), we did not calculate its velocity based on the transformed values and instead calculated the change in yaw angles prior to any coordinate transformations. This procedure preserves information on how a user rotates their head in the yaw axis, which body-space transformation loses by reorienting based on head yaw.
@@ 3.3.4. Dyadic Relationship Features
The next group of features describes the dynamics between the main user and reference user. Similar to prior work (Ishii et al., 2016b; Chen et al., 2024; Jokinen et al., 2013), we extracted the direct gaze angles from the main user to the reference user, and from the reference user to the main user. Specifically, we determined user A’s direct gaze angle towards user B by calculating the angle between user A’s forward head orientation in the yaw axis and the vector pointing from user A’s head position to user B’s head position in the horizontal plane. Similar to the egocentric motion features, we derived intermediate values (i.e., raw values, velocities) across the 1-second window and summarized them by their average, minimum, and maximum values. Given that users in the examined dataset were also free to move around the virtual environments, we further encoded interpersonal distance as the distance between two users’ head position in the horizontal plane.
One characteristic that direct gaze and interpersonal distance fails to capture is whether the two users are facing the same direction when they are not looking towards each other. For example, if both users are at a 90-degree direct gaze angle from the other user, they could either be facing the same direction or have their backs against one another. These two scenarios can suggest different dynamics as users facing the same direction and looking at similar parts of the environment may be more inclined to interact with each other. We therefore introduced a measurement of visual shared space to quantify how much of two users’ visual field of views overlap. Concretely, we first drew isosceles triangles with the vertex connecting the two equal sides of length located at the user’s head and oriented in the head’s forward direction in the horizontal plane. The angle between the two equal sides is set to the horizontal field of view of the Meta Quest 2 headset (i.e., 104 degrees). The raw value of the visual shared space between two users at distance is then defined as the amount of spatial overlap in () of the two triangles drawn with . In simpler terms, this measurement estimates how much mutual space the two users share in their field of views, assuming that their visual attention stretches a distance of . For our problem setup, we extracted values for at 1, 5, and 10 meters.
Using the raw values calculated for direct gaze, interpersonal distance and visual shared space, we again extracted the raw values as well as their changes (i.e., velocity) over the sampled windows and summarized the two measurements into their average, minimum, and maximum values.
@@ 3.3.5. Group Relationship Features
Finally, we encoded features between a user of interest (i.e., main or reference user) with the rest of the social group. Specifically, we extracted (1) the dyadic direct gaze angles from each user of interest to each remaining member of the group and (2) those from each member of the group to that user. Similar to the dyadic relationship features, we calculated the pairwise interpersonal distance between each of the two users with the rest of the group. Finally, we derived the visual shared space between each dyadic pair between a user of interest and the rest of the group at = 1, 5, 10 meters. From the raw values extracted for direct gaze, interpersonal distance, and visual shared space across the sampled window, we derived the average raw value and velocity for each dyadic pair, grouped them by whether they are related to the main or reference user, and finally summarized them by their average, minimum, and maximum values.
@@ 3.4. Turn-Taking Behavior Prediction Tasks
We are interested in understanding whether individual- and group- level differences and nonverbal and verbal behaviors can predict turn-taking behaviors (RQ1). Drawing from past literature on predicting speech behaviors (Ishii et al., 2016b, 2017; Chen et al., 2024; Lee et al., 2023; Parker, 1988), we focused on three prediction tasks, which at a high level captures the “what”, “who”, and “when” of turn-taking behaviors.
Turn Taking vs. Continuing Speech. The “what” poses the question of whether we can predict the type of turn-transition behavior. In particular, in line with past research (Ishii et al., 2016b; Lee et al., 2023; Jokinen et al., 2013), can we distinguish whether a turn-transition behavior was going to be a turn-transition to a new speaker (i.e., turn taking), or a continuing speech (i.e., turn keeping) from the previous speaker?
Next Speaker Prediction. The “who” focuses on predicting who the new speaker is prior to the start of a turn-taking behavior (i.e., clean turn taking, overlap turn taking). Predicting the next speaker is another common task used in modeling speech behavior (Ishii et al., 2016b; Parker, 1988; Ishii et al., 2016c, a).
@@ 3.5. Machine Learning Models
To study RQ1 and predict the three tasks outlined in Section 3.4, we began by extracting the features in the 1-second windows prior to the start of the turn-transition behaviors. For each task, we built and compared performance across four predictive models commonly used to predict individual characteristics and group dynamics (Cao et al., 2021; Miller et al., 2023b; Moore et al., 2021): logistic regression, multi-level perceptron (MLP) classifier, random forest classifier, and gradient boosting classifier. We formulated each task as a binary classification task and measured prediction performance using the area under the curve (AUC) of the receiver operating characteristics curve. We implemented all models using Python’s scikit-learn library (Fabian, 2011).
In evaluating each model, we employed cross-validation similar to (Frommel et al., 2020; Cao et al., 2021), and reported the averages and standard deviations of the AUC evaluated on the testing data across all folds. Once we partitioned the data into training and testing, we standardized all continuous features. For measuring accuracy, we reported the performance of models trained on 90% of the data and evaluated on the remaining 10% (i.e., k-fold 10). As a benchmark measurement, we partitioned the data such that the training data consisted of data points from 90% of the sessions and the testing was done on the remaining 10% (Sessioncv).
To investigate RQ2, the robustness of predictive models on unseen groups, activities, and weeks, we report two additional sets of results. First, we measured performance on unseen groups by partitioning the data such that training was performed on data collected from 90% of groups and testing was performed on samples collected from the remaining 10% (Groupcv). We then quantified performance across unseen weeks using models trained on all but one week’s data and tested on that remaining week (Weekcv). Finally, we report performance on the final week (i.e., week 4) when being trained on data collected over the first three weeks. As group activities differed each week, our evaluation metric for unseen activities is the same as that for unseen weeks.
@@ 4. Results
In this section, we report on the model performance for each of the three prediction tasks outlined in Section 3.4. Then, we present additional analyses to investigate feature importance and interpret how the features are related to turn-taking behaviors predictions.
@@ 4.1.1. Differentiating between Turn Taking and Continuing Speech
Random Forest Classifier | 0.76 (0.03) | 0.75 (0.03) | 0.76 (0.01) | 0.76 |
@@
We first built models for predicting whether a speech event was going to be a turn transition or one where the previous speaker will continue to speak. We chose to compare clean turn taking, and not overlap turn taking, to continuing speech as there are by definition pauses before both types of turn-transition behaviors. We defined positive samples as those corresponding to the beginning of each turn-taking behavior and negative samples as those corresponding to the start of continuing speech behaviors. We chose the main user as the upcoming speaker for the positive samples and a randomly selected user who was not the previous speaker for negative samples. For both positive and negative samples, the reference user was the previous speaker. After aggregating the data samples, we sampled the dataset such that there is an equal number of positive and negative data samples. This yielded a dataset with 15876 samples.
We present model performance in Table 2. Notably, we observe that the gradient boosting classifier outperformed the remaining three models across all four performance metrics, achieving a benchmark accuracy (i.e., Sessioncv) of 0.78 AUC, and 0.77–0.78 AUC across the three remaining performance metrics. Random forest classifiers achieved the second highest accuracy on the benchmark metric with a 0.76 AUC, followed by the MLP classifier at 0.72 AUC, and the logistic regression at 0.70 AUC. When evaluating on unseen groups, performance accuracies were slightly lower but generally robust across all four prediction models. Compared to the benchmark metric, the models achieve similar accuracies on unseen weeks and activities.
@@ 4.1.2. Predicting the Next Speaker
MLP Classifier | 0.72 (0.04) | 0.71 (0.04) | 0.71 (0.01) | 0.73 |
Random Forest Classifier | 0.75 (0.03) | 0.74 (0.04) | 0.75 (0.01) | 0.76 |
@@
In the second task, we predicted the upcoming speaker of a turn transition event. Specifically, we aggregated positive samples by extracting those corresponding to moments at the beginning of turn-taking events (i.e., clean turn taking, overlap turn taking), where the main user is set to the upcoming speaker. We then collected negative samples at the same moments, but with the main speaker set to a user who is neither the upcoming nor the previous speaker. The reference user is set to the previous speaker for all samples. Upon collecting all positive and negative samples, we down-sampled the negative labels such that there is an equal number of positive and negative data samples. The final dataset consisted of 22316 samples.
Table 3 summarizes our results for predicting the next speaker. Gradient boosting classifier achieved the highest accuracies across all four performance metrics, with an AUC of 0.77 on the benchmark metric, followed by the random forest classifier at 0.75 AUC, the logistic regression at 0.73 AUC, and finally the MLP classifier at 0.72 AUC. All four prediction models were robust when tested on unseen groups and weeks, with the performance on unseen groups again being lower but comparable to the benchmark evaluations.
@@ 4.1.3. Predicting the Timing of Turn Taking
Finally, we were interested in understanding whether we can differentiate between moments associated with an upcoming speaker right before a turn transition event and those sampled prior to these moments. We therefore collected positive samples by extracting samples corresponding to moments before the start of turn-taking speech events (i.e., clean turn taking, overlap turn taking), with the main user set to the upcoming speaker. For negative samples, we aggregated samples corresponding to moments that were prior to the 1-second window immediately before the start of turn transition event. To do this, we began by sampling moments at windows starting 2, 4, 6, 8, 10, and 12 seconds prior to the start of turn transition events. We then filtered out data samples for which any speech event had occurred between the earlier moment and the start of the turn transition event for the upcoming speaker. In other words, the negative samples consisted only of moments when the upcoming speaker is not preparing to speak. This sampling technique is similar to Chen et al. (Chen et al., 2024) in that we sampled data points by employing a sliding window. For both positive and negative samples, the reference user is set to the previous speaker at the sampled moment. After aggregating the samples, we again sampled the data so that there is an equal number of positive and negative samples. The dataset for this task contained 22316 samples.
We detail performance results in Table 4. Compared with the performance on the previous two tasks, the accuracies were considerably lower. The best performing model, the gradient boosting classifier, achieved an accuracy of 0.72 AUC on the benchmark metric, while the logistic regression model recorded the lowest accuracy of 0.61 AUC on the same metric. The random forest and MLP classifiers had benchmark accuracies of 0.68 and 0.63 AUC, respectively. Performance on unseen groups, activities, and weeks were robust, with models reporting slightly lower but similar accuracies on the cross validation evaluations with unseen groups.
@@ 4.2. Feature Importance Analysis
![Refer to caption](extracted/5634250/figures/feature-importance-v6.png)
@@
To understand how features are related to model performance and predictions RQ3, we begin by computing the mean decrease in accuracy (MDA), a technique used for probing feature importance in machine learning models (Miller et al., 2020; Han et al., 2016; Nicodemus, 2011). Specifically, we first organized features based on their feature groups and definition, for example splitting up egocentric motion features by the user, tracked points and axis of motion. For each task, we built models used to compute the benchmark cross validation performance (i.e., Sessioncv) and evaluated the models on each testing set with a particular set of features randomly shuffled. This procedure breaks up any relationship the set of features may have with the model prediction. We then calculated the average change in the testing AUC across all folds and present them in Figure 3. Note that MDA is defined as the amount of decrease in accuracy after feature shuffling, meaning that lower values (i.e., greater decrease in accuracy) in Figure 3 represents higher MDA. We omitted plotting egocentric features with moderate importance, but included all features in Appendix B.
@@
Analysis of speech-related features showed the highest importance on the user’s previous turn time, followed by prior speech sequence, user turn count, and whether or not the user has spoken earlier in the current session. Notably, the main user’s previous turn time feature, the feature importance for the turn taking vs. continuing speech task was the highest, followed by the timing prediction and the next speaker prediction tasks. For features on the main user’s turn count and prior speech sequence, models built to predict the next speaker exhibited the highest importance, followed by those for the turn taking vs. continuing speech task and lastly the timing prediction task. Our results on the individual and group differences feature revealed a comparably high importance of the main user personality, followed by group size, group personality, and lastly the reference user personality.
@@
For egocentric motion features, the top six features were those related to the main user’s and reference user’s (1) head yaw rotation, (2) head y-axis position, and (3) left hand y-axis position. Within the top six features, the feature importance for those related to the main user were in most cases higher for the timing prediction task compared to the other tasks. For dyad relationship features, those related to direct gaze had greater importance than those related to the visual shared space and interpersonal distance, though their change in AUC near zero implies low feature importance. In terms of group relationship features, we found that the features related to the group’s direct gaze towards the reference and main users exhibited greater importance than the remaining features on visual shared space, interpersonal distance, and direct gaze from the main and reference user to the rest of the group.

### v2

@@ 3.3.3. Egocentric Motion Features
Egocentric motion features capture the pose and motion of individuals. For this, we extracted egocentric motion features from both the main and reference users. For each user, we extracted the position (i.e., x, y, z) and orientation (i.e., pitch, yaw, roll) of a user’s headset and two controllers by both their raw values and velocities. From the raw values and velocities across the 1-second window, we summarized each of the 6 degrees of freedom by their average, minimum, and maximum. We transformed the features corresponding to the raw values into the body-space coordinate system proposed in Miller et al. (Miller et al., 2023b) and centered the horizontal coordinates (i.e., x, z) such that the individual’s head position is at the origin of the horizontal plane, as opposed to their global coordinates in the virtual space. The rationale behind this transformation is to encode the egocentric poses by mapping the raw coordinates and orientations to a coordinate system based on the individual’s forward head direction. This allows the features to encode poses even when individuals move around physical or virtual environments.
When calculating the velocity of positions and orientations, we derived velocities from raw values after transforming the original values into the body-space coordinates without the centering operation. Not centering the coordinates based on the head position allows the features to retain information of the user’s head position velocities in the horizontal plane. When encoding head yaw, we did not calculate its velocity based on the transformed values but instead on the yaw angles prior to any transformations. This procedure preserves information on how a user rotates their head in the yaw axis, which body-space transformation loses by reorienting based on head yaw.
@@ 3.3.4. Dyadic Relationship Features
The next feature group describes the dynamics between the main and reference users. Similar to prior work (Ishii et al., 2016b; Chen et al., 2024; Jokinen et al., 2013), we extracted the direct visual attention (i.e., gaze direction approximated using head orientation) from the main user to the reference user, and vice versa. Specifically, we determined user A’s direct visual attention towards user B by calculating the angle between user A’s forward head orientation in the yaw axis and the vector pointing from user A’s head position to user B’s head position in the horizontal plane. Similar to the egocentric motion features, we derived the raw values and velocities across the 1-second window and summarized them by their average, minimum, and maximum. Since users moved around virtual environments in the dataset, we also encoded interpersonal distance as the distance between two users’ head position in the horizontal plane.
One characteristic that direct visual attention fails to capture is whether two users are facing the same direction. For example, if both users’ head orientations are at 90 degrees from the other user, they could be facing the same direction or have their backs against one another. These two scenarios can entail different dynamics as users facing the same direction are likely looking at similar parts of the environment. We therefore introduced a measurement of visual shared space quantifying how much users’ visual field of views overlap. Concretely, we first drew isosceles triangles with the vertex connecting the two equal sides of length located at the user’s head and oriented in the head’s forward direction in the horizontal plane. The angle between the two sides is set to the horizontal field of view of the Meta Quest 2 headset (104 degrees). The raw value of the visual shared space between two users at distance is defined as the amount of spatial overlap in of the two triangles drawn with . In simpler terms, this measurement estimates how much virtual space the two users share in their field of views, assuming that their visual attention stretches meters. For our setup, we extracted values for at 1, 5, and 10 meters.
Using the raw values calculated for direct visual attention, interpersonal distance and visual shared space, we extracted the raw values and their changes (i.e., velocity) over the sampled windows and summarized the two measurements into their average, minimum, and maximum.
@@ 3.3.5. Group Relationship Features
Finally, we encoded features between a user of interest (i.e., main or reference user) with the group. Specifically, we extracted (1) the dyadic direct visual attention angle from each user of interest to each group member and (2) those from each group member to that user. Similar to the dyadic relationship features, we calculated the interpersonal distance between each of the two users with the rest of the group. Finally, we derived the visual shared space between each dyadic pair between a user of interest and the rest of the group at = 1, 5, 10 meters. From the raw values extracted for direct visual attention, interpersonal distance, and visual shared space across the sampled window, we derived the average value and velocity for each dyadic pair, grouped them by whether they are related to the main or reference user, and finally summarized them by their average, minimum, and maximum.
@@ 3.4. Turn-Taking Behavior Prediction Tasks
Drawing from past literature on predicting speech behaviors (Ishii et al., 2016b, 2017; Chen et al., 2024; Lee et al., 2023; Parker, 1988), we focused on three prediction tasks, which capture the “what”, “who”, and “when” of turn-taking behaviors.
Turn Taking vs. Continuing Speech. The “what” poses the question of whether we can predict the type of turn-transition behavior. In particular, in line with past research (Ishii et al., 2016b; Lee et al., 2023; Jokinen et al., 2013), can we distinguish whether a turn-transition behavior will be a turn-transition to a new speaker (i.e., turn taking), or a continuing speech from the previous speaker (i.e., turn keeping)?
Next Speaker Prediction. The “who” focuses on predicting who the new speaker is prior to the start of a turn-taking behavior (i.e., clean turn taking, overlap turn taking). Next speaker prediction is another common task used in modeling speech behavior (Ishii et al., 2016b; Parker, 1988; Ishii et al., 2016c, a).
@@ 3.5. Machine Learning Models
To study RQ1, we predicted the three tasks outlined in Section 3.4 by first extracting the features in the 1-second windows prior to the start of the turn-transition behaviors. For each task, we built and compared performance across four predictive models commonly used to predict individual characteristics and group dynamics (Cao et al., 2021; Miller et al., 2023b; Moore et al., 2021): logistic regression, multi-level perceptron (MLP) classifier, random forest classifier, and gradient boosting classifier. We formulated each task as a binary classification task and measured prediction performance using the area under the curve (AUC) of the receiver operating characteristics curve. We implemented all models using Python’s scikit-learn library (Fabian, 2011).
For evaluation, we employed cross-validation similar to (Frommel et al., 2020; Cao et al., 2021)333We conducted an additional analysis that focused on evaluating model performance across sessions, groups, and weeks. Since the results largely align with our cross-validation evaluations, we include this analysis in Appendix D.. To start, we partitioned the data into training and testing and standardized all continuous features. For measuring accuracy, we reported the performance of models trained on 90% of the data and evaluated on the remaining 10% (i.e., k-fold 10). Specifically, we calculated the averages and standard errors of the AUC evaluated on the testing data across all folds. As a benchmark measurement, we partitioned the data such that we trained on data from 90% of the sessions and tested was on the remaining 10% (Sessioncv).
We report two additional sets of results evaluating the robustness of models on unseen groups, activities, and weeks (RQ2). First, we measured performance on unseen groups by partitioning the data such that we trained on data from 90% of groups and tested on the remaining 10% (Groupcv). We then quantified performance across unseen weeks using models trained on all but one week’s data and tested on the remaining week (Weekcv). Finally, we report performance on the final week after training models on data from the first three weeks. As group activities differed weekly, our evaluation metric for unseen activities is the same as that for unseen weeks.
@@ 4. Results
In this section, we report on model performance (RQ1) and their robustness (RQ2) for the prediction tasks outlined in Section 3.4. Then, we present analyses investigating feature importance and how the features are related to model predictions (RQ3).
@@ 4.1.1. Differentiating between Turn Taking and Continuing Speech
Random Forest Classifier | 0.77 (0.00) | 0.75 (0.01) | 0.76 (0.00) | 0.76 |
@@
We built models for predicting whether a speech event was going to be a turn transition or one where the previous speaker will continue to speak. We compared clean turn taking, and not overlap turn taking, to continuing speech as there are by definition pauses before both types of turn-transition behaviors. We defined positive samples as those corresponding to the beginning of turn-taking behaviors and negative samples as those corresponding to the start of continuing speech behaviors. We chose the upcoming speaker as a main user for positive samples and a randomly selected user who is not the previous speaker for negative samples. For both positive and negative samples, the reference user was the previous speaker. We randomly sampled the dataset to maintain an equal number of positive and negative samples, which yielded 15844 samples.
We present model performance in Table 3. Notably, gradient boosting classifiers outperformed other models across all performance metrics, achieving a benchmark accuracy (i.e., Sessioncv) of 0.78 AUC, and 0.77–0.78 AUC across the three remaining metrics. Random forest classifiers achieved the second highest accuracy on the benchmark metric with a 0.77 AUC, followed by the MLP classifier at 0.72 AUC, and the logistic regression at 0.71 AUC. When evaluating on unseen groups, performance accuracies were slightly lower but generally robust across all four models. Compared to the benchmark metric, models achieve similar accuracies on unseen weeks and activities.
@@ 4.1.2. Predicting the Next Speaker
MLP Classifier | 0.72 (0.01) | 0.71 (0.01) | 0.72 (0.01) | 0.74 |
Random Forest Classifier | 0.75 (0.01) | 0.74 (0.01) | 0.74 (0.00) | 0.75 |
@@
We predicted the upcoming speaker of a turn-taking event. Specifically, we aggregated positive samples by extracting those corresponding to moments at the beginning of turn-taking events (i.e., clean turn taking, overlap turn taking), where the main user is set to the upcoming speaker. We then collected negative samples at the same moments, but with the main speaker set to a user who is neither the upcoming nor the previous speaker. The reference user is set to the previous speaker for all samples. We randomly resampled negative labels to balance positive and negative samples, yielding 22316 samples.
Table 4 summarizes our results. Gradient boosting classifier achieved the highest accuracies across all four performance metrics, with an AUC of 0.77 on the benchmark metric, followed by the random forest classifier at 0.75 AUC, the logistic regression at 0.73 AUC, and finally the MLP classifier at 0.72 AUC. All four prediction models were robust when tested on unseen groups and weeks, with the performance on unseen groups again being lower but comparable to the benchmark evaluations.
@@ 4.1.3. Predicting the Timing of Turn Taking
Finally, we aimed to understand whether we can differentiate between moments associated with an upcoming speaker right before a turn taking event and those sampled prior to these moments. We first collected positive samples by extracting samples corresponding to moments before the start of turn-taking events (i.e., clean turn taking, overlap turn taking), with the main user set to the upcoming speaker. For negative samples, we aggregated samples corresponding to moments that were prior to the 1-second window immediately before the start of turn-taking event. For this, we began by sampling moments at windows starting 2, 4, 6, 8, 10, and 12 seconds prior to the start of turn transition events. We then filtered out data samples for which any speech event had occurred between the earlier moment and the start of the turn transition event for the upcoming speaker. In other words, the negative samples consisted only of moments when the upcoming speaker is not preparing to speak. This sampling technique is similar to Chen et al. (Chen et al., 2024) as we sampled through a sliding window. For both positive and negative samples, the reference user is set to the previous speaker at the sampled moment. We randomly resampled the aggregated data to ensure an equal number of positive and negative samples, resulting in 22316 samples.
We detail performance results in Table 5, which were lower than those from previous tasks. The best performing model, the gradient boosting classifier, achieved an accuracy of 0.72 AUC on the benchmark metric, while the logistic regression model recorded the lowest accuracy of 0.61 AUC on the same metric. The random forest and MLP classifiers had benchmark accuracies of 0.68 and 0.63 AUC, respectively. Performance on unseen groups, activities, and weeks were robust, with models reporting slightly lower but similarly accuracies on other cross validation evaluations.
@@ 4.2. Feature Importance Analysis
4.2.1. Mean Decrease in Accuracy Based on Gradient Boosting Classifiers
@@
To understand how features are related to model performance and predictions (RQ3), we first computed the mean decrease in accuracy (MDA) for the top performing model (i.e., gradient boosting classifier), a technique used for probing feature importance in machine learning models (Miller et al., 2020; Han et al., 2016; Nicodemus, 2011). Specifically, we organized features based on their feature groups and definition, for example splitting up egocentric motion features by the user, tracked points, and axis of motion. For each task, we built models used to compute the benchmark cross validation performance (i.e., Sessioncv) and evaluated them on each testing set with a particular set of features randomly shuffled. This procedure breaks up any relationship the set of features may have with the model prediction. We then calculated the average change in the testing AUC across all folds and present them in Figure 3. We omitted plotting egocentric features with moderate importance, but included all features in Appendix B.
@@
Analysis of speech-related features showed the highest importance on the user’s previous turn time, followed by prior speech sequence, user turn count, and whether the user has spoken earlier in the session. For the main user’s previous turn time feature, feature importance for the turn taking vs. continuing speech task was the highest, followed by the timing prediction and next speaker prediction tasks. For features on the main user’s turn count and prior speech sequence, they exhibited the highest importance for models built to predict the next speaker, followed by the turn taking vs. continuing speech task, andlastly the timing prediction task. Our results on the individual and group differences features revealed a comparably high importance of the main user personality, followed by group size, group personality, and lastly the reference user personality.
@@
For egocentric motion features, the top six features were related to the main and reference users’ head yaw rotation, head y-axis position, and left hand y-axis position. Within the top six features, feature importance for those related to the main user were in most cases higher for the timing prediction task compared to other tasks. For dyad relationship features, those related to direct visual attention had greater importance than those related to the visual shared space and interpersonal distance, though their change in AUC near zero implies low importance. In terms of group relationship features, features related to the group’s direct visual attention towards the reference and main users exhibited greater importance than the remaining features.
4.2.2. Feature Significance Based on Logistic Regression Models
We investigated predictor significance of feature groups using logistic regressions. Benefits of studying linear models include their high interpretability and simplified independence assumption affording well-defined hypothesis tests for evaluating feature significance (Couronné et al., 2018). In contrast, recent methods for interpreting more complex models such as the random forest and gradient boosting classifiers have not addressed evaluations of feature importance through statistical tests (Greenwell et al., 2018; Lundberg, 2017; Lundberg et al., 2018). For this, we grouped features based on their constructs. Then, for each trained logistic regression, we conducted Wald tests to evaluate joint linear hypotheses, specifically testing whether all features related to a given construct significantly contribute to the model. We present the full results in Appendix C but highlight general trends for the remainder of this section. We evaluated significance at =.05.
Specifically, all but one group of speech-related features significantly contributed to model prediction across all three tasks (s¡.046). For individual and group differences features (i.e., personality, group size), most significantly impacted the tasks of predicting turn taking vs. continuing speech and next speaker prediction (s¡.020). In contrast, for individual and group differences features, only features related to the previous speaker’s personality (¡.001) significantly predicted turn-taking timing. For dyad-related features, visual attention features between the reference and main users significantly predicted the next speaker (s¡.001) while the interpersonal distance between the reference and main users was significant in distinguishing between turn taking and continuing speech behaviors (=.004). For features related to groups dynamics, all constructs significantly predicted turn taking vs. continuing speech (s¡.017). Group features related to visual attention were also significant in predicting at least one of the two remaining tasks. Finally, for egocentric features, 18 groups of constructs were significant in predicting turn taking vs. continuing speech (s¡.035), while only 6 were significant in predicting the next speaker (s¡.027). Notably, all 6 were related to the main user’s egocentric behavior. For predicting the timing of turn taking, 12 groups of constructs were significant (s¡.043).